OpenStudy (anonymous):

Use common denominators to determine whether the ratios form a proportion. 6 : 9 and 9 : 12 A. Yes. They form a proportion. B. No. They do not form a proportion.

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

they both can go into each other

OpenStudy (destinymasha):

tell me how...

OpenStudy (destinymasha):

[i didnt say u were wrong]

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i not good at math

OpenStudy (anonymous):

:(

OpenStudy (destinymasha):

i don't think its a proportion.....

OpenStudy (destinymasha):

@Tellme2242

OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 12 : 44 and 3 : 11 A. No. They do not form a proportion. B. Yes. They form a proportion.

OpenStudy (destinymasha):

yes--- 44 divided by 4 = 11 , and 12 divided by 4 = 3

OpenStudy (anonymous):

Which ratio can be used to form a proportion with the given ratio? 7 to 12 A. 8 to 13 B. 10 to 15 C. 14 to 19 D. 21 to 36

OpenStudy (destinymasha):

7/12 --- 7 x 3 = 21, and 12x 3 = 36 , so 7/12 = 21/36

OpenStudy (anonymous):

There are 4 boys and 6 girls in a singing group. The ratio of boys to girls on the debate team is proportional to the ratio of boys to girls in the singing group. Which could be the number of boys and girls on the debate team? A. 6 boys and 8 girls B. 6 boys and 9 girls C. 10 boys and 12 girls D. 16 boys and 36 girls

OpenStudy (anonymous):

is it c?

OpenStudy (destinymasha):

ITS D! ITS D!

OpenStudy (destinymasha):

the answer is D!!!

OpenStudy (destinymasha):

4x 4 = 16, and 6 x 6 = 36